# supplementary_materials

### v1

@@ IV.2 Critical fluctuations mediated pair-density wave
The fluctuations about the quantum critical point are modelled by two bosonic modes $\delta\chi_{\mathbf{r}}=b_{\pm}(\mathbf{r})$, which condense at the translation-symmetry breaking exciton condensate momenta ±$\pm\mathbf{Q}$. From the $C_{2z}$ symmetry (that relates $+\mathbf{Q}$ to $-\mathbf{Q}$ condensates), the interaction of the bosons with the fermionic (electon/hole) excitations is captured by the Hamiltonian, $H=H_{f}+H_{b}+H_{f-b}$, where
@@
Here $H_{f}$ is the bare action of the fermions that form electron and hole Fermi surfaces. The density wave operator is $\rho_{DW}(\mathbf{q})=\int_{\mathbf{k}}f^{\dagger}_{v,\mathbf{k}+\mathbf{q}}f_{c,\mathbf{k}}$, the mass (speed) of the bosons is given by $m_{b}$ ($c_{b}=1$, for simplicity), $\kappa$ captures the anisotropy in the boson dispersion, and we have only kept the leading momentum independent coupling constant $g$. Our approximation of considering only static fluctuations is justified in the presence of a sufficiently large $m_{b}$. Integrating out the bosons we arrive at an effective interaction term between the fermions, $-\frac{1}{A}\sum_{\mathbf{q}}V_{DW}(\mathbf{q})\rho_{DW}(\mathbf{q},\Omega=0)\rho^{\dagger}_{DW}(\mathbf{q},\Omega=0)$, where
$V_{DW}(\mathbf{q})=g^{2}\sum_{s=\pm}\frac{1}{(q_{x}-sQ)^{2}+\kappa^{2}q_{y}^{2}+m_{b}^{2}}.$ (8)
We estimate $g$, $\kappa$ and $m_{b}$ as a function of $n$ and $V_{b}$ by using short-ranged (Hubbard) bare interactions required to approximately reproduce the quantum critical line in Fig. 2 (see Appendix C). Since $\rho_{DW}$ is an interlayer density operator, the superconducting instabilities due to the interaction term occurs in the interlayer channel, i.e., the Cooper pairs are formed out of electrons that reside in opposite layers. The non-vanishing superconducting instability is found to be a pair-density wave (PDW) state with pairing momentum $\mathbf{Q}$: $\langle f_{v,\mathbf{k}+\mathbf{Q}}f_{c,-\mathbf{k}}\rangle,\langle f_{c,\mathbf{k}+\mathbf{Q}}f_{v,-\mathbf{k}}\rangle\neq 0$ (as depicted in Fig. 5), as opposed to the standard zero-momentum pairing. Since the PDW introduces a periodicity along the $x$-direction in the Bogoliubov-de Gennes (BdG) Hamiltonian, the momentum $k_{x}$ is defined in the Brillouin zone: $k_{x}\in[-Q,Q)$. We correspondingly fold the states $f_{c,\mathbf{k}+2m\mathbf{Q}}$, $f^{\dagger}_{v,-\mathbf{k}+(2m+1)\mathbf{Q}}$ for integer $m$ onto the first Brillouin zone, and the mean-field decoupled terms in the pairing channel are of the form,
$\sum_{\mathbf{k},m}$ $\Delta_{+,m}(\mathbf{k})f^{\dagger}_{c,\mathbf{k}+2m\mathbf{Q}}f^{\dagger}_{v,-\mathbf{k}-(2m-1)\mathbf{Q}}$
[…]
where $\Delta_{\pm}(\mathbf{k})$ are superconducting gap equations with pairing momentum ±$\pm\mathbf{Q}$, and $m\in\mathbb{Z}$. For our calculations, we truncate to keep the five lowest energy bands and self-consistently obtain $\Delta_{+}=\Delta_{-}$ by solving the BdG Hamiltonian (refer to Appendix E for the details). Figure 5 (top) depicts the strength of the PDW order parameter for ordering at ±$\pm\mathbf{Q}$: $\frac{1}{N_{k}}\sum_{\mathbf{k},m}\langle f_{c,\mathbf{k}+2m\mathbf{Q}}f_{v,-\mathbf{k}-(2m\pm 1)\mathbf{Q}}\rangle$ as a function of $n$ and $V_{b}$ (we obtain the same strength for $\mathbf{Q}$ and $-\mathbf{Q}$). As seen, the PDW superconductor is stabilized near the quantum critical line, where the interactions mediated by the soft bosons $b_{\pm}$ are strongest. The development of the PDW state requires a critical interaction strength, therefore, for small values of $g$ or large values of $m_{b}$, superconductivity disappears, as seen in Fig. 5. We find that the obtained PDW gap functions have an anisotropic $p$-wave character (see Fig. 5, bottom panel) with nodes at $k_{x}=0,-Q$. The resulting BdG spectrum consists of gapless quasiparticles that form a Fermi surface, as is typical for PDW superconductors Agterberg et al. (2020).
@@
Figure 5: Critical fluctuation-mediated pair-density wave superconductivity. Top: Order parameter of the pair-density wave (PDW) superconductor plotted on the electron-hole plasma side of the phase diagram (lower triangle). Bottom: The PDW gap function for $n=-0.36\times 10^{12}$ cm-2 and $V_{b}=4.2$ meV plotted in the first Brillouin zone. Dashed lines are Fermi surfaces of $f_{c,\mathbf{k}}$, $f_{v,\mathbf{k}+\mathbf{Q}}$ and $f_{c,\mathbf{k}-\mathbf{Q}}$. Results were obtained using a $79\times 79$ momentum mesh.
V Experimental Signatures
@@
Having demonstrated that bilayer semiconductor platforms are promising for the realization of exciton density wave phases – and foment superconducting instabilities within the phase or near the phase transition – we now focus on the question of what probes can distinctly identify the various phases and phenomena discussed in this work. The X-DW order as well as both the superconducting phases involve an inter-layer condensate (the Goldstone mediated superconductor also involves intra-layer pairing). This suggests drag transport – with separate electrical contacts, current is driven in one layer and the voltage generated is measured in the opposite layer – as a natural probe for these phenomena Eisenstein (2014). Analogous to the exciton condensates found in bilayer quantum Hall systems, the drag (and counterflow) resistivity will be zero in the X-DW phase even though the excitons condense at a finite momentum. We direct the reader to Appendix G for arguments based on effective response actions. Since the PDW superconductor is an interlayer Cooper pair condensate, the drag resistivity is zero even in this case. However in contrast to the X-DW, the PDW state will exhibit superconductivity when the layers are contacted together. Finally, in the quantum critical NFL regime, the hotspot contribution to AC counterflow transport could potentially contain signatures of the quantum criticality, if they can be isolated from the contributions from the other “cold” parts of the Fermi surface Shi et al. (2024).
Though our mean-field analysis suggest the formation of collinear X-DW over spiral X-DW, it is instructive to determine whether this may be experimentally verified. We recall that both the collinear and the spiral X-DW phases show the same qualitative behavior in drag transport. However, the presence of intra-layer charge density modulation in the case of the collinear X-DW can distinguish the two possibilities, for example by scanning tunneling microscopy. Another distinguishing feature is the $C_{2z}$ symmetry breaking in the case of the spiral X-DW, which could be detected in angle-resolved transport experiments Wu et al. (2020).
Finally, we discuss the response of the system to an in-plane magnetic field. The orbital coupling of the magnetic field will cause the two-dimensional momenta of the electrons in the two layers to shift in opposite directions, explicitly breaking $C_{2z}$. (For magnetic field along $\hat{x}$, the vector potential in the Landau gauge, $\mathbf{A}\sim z\hat{y}$, which is opposite for the two layers located at $z\pm d/2$.) The equally susceptible wave-vectors for exciton condensation will then be $\mathbf{Q}_{1}$ and $-\mathbf{Q}_{2}$ for $\mathbf{Q}_{1}\neq\mathbf{Q}_{2}$. In this case, we expect the collinear X-DW to get modified to a quasi-periodic ordered state where the strength of the exciton order parameter varies with a wavevector $(\mathbf{Q}_{1}+\mathbf{Q}_{2})/2$ and the phase of the order parameter varies with $(\mathbf{Q}_{1}-\mathbf{Q}_{2})/2$.
VI Outlook
@@
We close with a discussion of promising future research directions related to our work. We have presented two forms of superconductivity – one arising due to Goldstone fluctuations and another due to quantum critical fluctuations – that can be realized at zero temperature. A natural direction to explore would be finite temperature energy competition between the two superconductors (for instance, the critical temperature $T_{c}$ of each and the dependence on $n,V_{b}$). Studying the energetic competition between the two superconductors and potential phase transitions between them could also be an avenue for future exploration.
@@
Recent experiments have presented evidence for the existence of a Fermi liquid state of trions (bound state of a fermion and an exciton) Nguyen et al. (2023). It would be interesting to incorporate trions in our framework, which would provide an alternate route for the excitons to become gapped by starting from the exciton condensed phase (as studied in the cold-atom context in Ref. Altman et al. (2012)). In particular, some of the condensed excitons could bind with an electron or a hole to form a trion Fermi sea state and effectively withdraw from participating in the condensate. If the trion Fermi sea becomes sufficiently large, it could completely gap out the excitons. Understanding the energetic competition of this alternate instability, in relation to the aforediscussed superconductivity, would be an interesting direction for future work.
@@
Finally, the possibility for exciton driven correlated phenomena has also been observed in compensated twisted double bilayer graphene Ghosh et al. (2022), where the formation of a finite-momentum exciton condensate heralds non-Fermi liquid instabilities about the transition Ghorai et al. (2023). It would be intriguing to examine whether unconventional superconductivity may also triggered in this system.
VII Acknowledgements
@@
Appendix A Microscopic Parameters of Model
In this study, we take WS2/WTe2 as a representative material in our numerics. This heterostructure is expected to have a type-II band alignment with the conduction (valence) band belonging to the WS2 (WTe2) layer. Let us define a microscopic lattice length scale $a=0.32\text{nm}$. This is not necessarily the lattice constant of the semiconductors under consideration, but a convenient length scale that we use to express other physical parameters. The effective valence and conduction masses are $m_{v,x}a^{2}=0.89\text{eV}^{-1}$, $m_{v,y}a^{2}=0.54\text{eV}^{-1}$v et al. (2015), $m_{c,x}a^{2}=m_{c,y}a^{2}=0.33\text{eV}^{-1}$ Conti et al. (2020), with a gate distance of $d_{g}=1.5\text{nm}$ and interlayer distance $d=3.5$ nm. For the Hubbard approximation models, we adopt a rectangular momentum grid of cutoff length $k_{c,x}=0.1a^{-1}$ along the $x$-direction, and $k_{c,y}=k_{c,x}\left(\frac{m_{v,y}m_{c,y}}{m_{v,x}m_{c,x}}\right)^{1/4}$ along the $y$-direction. The chemical potential and bias voltage defined in the main text determine the Fermi momenta $k_{F,c/v,x/y}$ of the carriers in the conduction/valence ($c$/$v$) layers along the $\hat{x}/\hat{y}$ directions.
@@ Appendix A Microscopic Parameters of Model
Appendix B Density Wave Instability
At the non-interacting level, the system is described by an annualar Fermi surface, with an inner electron-like Fermi surface and an outer hole-like Fermi surface. Previous studies Pieri et al. (2007); Varley and Lee (2016) have established an instability of the annular Fermi surface to an interlayer charge density wave, which can be described by the order parameter $\langle f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}f_{c,\mathbf{k}}\rangle$. We obtain the conditions required for such an instability for our model within mean-field theory. Performing a mean-field decoupling in the spiral X-DW channel of interest, we obtain the following mean-field Hamiltonian:
@@ Appendix B Density Wave Instability
$\Sigma$ $=-\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}U(\mathbf{k}^{\prime}-\mathbf{k})f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{Q}}\langle f^{\dagger}_{v,\mathbf{k}^{\prime}+\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}\rangle+h.c.$ (11)
@@
$H_{\text{mf}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}}&f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}}&-\chi_{\mathbf{k}}\\ -\chi^{*}_{\mathbf{k}}&\xi_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}}\\ f_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}$ (12)
@@
$\chi_{\mathbf{k}}=\frac{1}{A}\sum_{\mathbf{k}^{\prime}}V(\mathbf{k}-\mathbf{k}^{\prime})\left(n_{F}(\xi_{\mathbf{k}^{\prime}}-E_{\mathbf{k}^{\prime}})-n_{F}(\xi_{\mathbf{k}^{\prime}}+E_{\mathbf{k}^{\prime}})\right)\frac{\chi_{\mathbf{k}^{\prime}}}{2E_{\mathbf{k}^{\prime}}}$ (13)
@@
This formulation to determine the exciton order parameter considers a select number of momentum shells/bands. For the exciton density wave instabilities (both spiral and collinear X-DW), we hence also perform fully self-consistent Hartree-Fock theory, including up to 7 $\times$2 bands; the number of bands selected is such that the UV cutoff is reached. We direct the reader to Ref. Dong et al. (2023) for analogous Hartree-Fock calculations performed in rhombohedral stacked multilayer graphene. We choose a UV cutoff of ±$\pm 0.1/a$ in the $\hat{y}$-direction, and the first Brillouin zone along the $\hat{x}$-direction bounded by the ±$\pm|{\mathbf{Q}}|/2$. The dielectric constant is taken to be $\epsilon=10$; typical choices of the hBN dielectric environment is $5\sim 10$ to account for additional screening by the mobile charge carriers in a system with finite carrier density. As described in the main text, our focus is on the inter-layer instabilities, rather than the intra-layer instabilities. This amounts to suppressing the intra-layer Hartree and Fock contributions in the decoupling.
We present in Fig. 6 a representative electronic bandstructure depicting the free (black) and finite exciton density wave order parameter (red). As seen, a electronic bands still cross the Fermi level, despite a finite exciton condensate, consistent with the three-band Hubbard approximation presented in Fig. 3 in Sec. III in the main text.
Figure 6: Self-consistent electronic bandstructure at $k_{x}=-0.00778$ for $n=-0.26\times 10^{12}$ cm-2 and $V_{b}=8.6$ meV. The red/black colors indicate the bandstructure in the presence/absence of the collinear exciton condensate. Momentum mesh size of 15$\times$15, and up-to 7$\times$2 (including both conduction and valence) folded bands was used.
Appendix C Estimation of $V_{DW}$

### v2

@@ Abstract
The fluctuations about the quantum critical point are modelled by two bosonic modes $\delta\chi_{\mathbf{r}}=b_{\pm}(\mathbf{r})$, which condense at the translation-symmetry breaking exciton condensate momenta $\pm\mathbf{Q}$. From the $C^{\prime}_{2z}$ symmetry (that relates the $\pm\mathbf{Q}$ condensates), the interaction of the bosons with the fermionic (electron/hole) excitations is captured by the Hamiltonian, $H=H_{f}+H_{b}+H_{f-b}$, where
@@
Here $H_{f}$ is the bare action of the fermions that form electron and hole Fermi surfaces. The density wave operator is $\rho_{DW}(\mathbf{q})=\int_{\mathbf{k}}f^{\dagger}_{v,\mathbf{k}+\mathbf{q}}f_{c,\mathbf{k}}$, the mass (speed) of the bosons is given by $m_{b}$ ($c_{b}=1$, for simplicity), $\kappa$ captures the anisotropy in the boson dispersion, and we have only kept the leading momentum independent coupling constant $g$. Our approximation of considering only static fluctuations is justified in the presence of a sufficiently large $m_{b}$. Integrating out the bosons we arrive at an effective interaction term between the fermions, $\frac{1}{A}\sum_{\mathbf{q}}V_{DW}(\mathbf{q})\rho_{DW}(\mathbf{q},\Omega=0)\rho^{\dagger}_{DW}(\mathbf{q},\Omega=0)$, where
$V_{DW}(\mathbf{q})=g^{2}\sum_{s=\pm}\frac{1}{(q_{x}-sQ)^{2}+\kappa^{2}q_{y}^{2}+m_{b}^{2}},$ (5)
where the constants $g$, $\kappa$ and $m_{b}$ are estimated (as functions of $n$ and $V_{b}$) by using a Hubbard interaction that approximately reproduces the quantum critical line in Fig. 2 (see [1]).
@@
Figure 5: Critical fluctuation-mediated pair-density wave superconductivity. Top: Order parameter of the pair-density wave (PDW) superconductor plotted on the electron-hole plasma side of the phase diagram (lower triangle). Bottom: The PDW gap function for pairing momentum $\mathbf{Q}$ plotted in the first Brillouin zone. Parameters: $n=-0.36\times 10^{12}$ cm-2 and $V_{b}=4.2$ meV. Dashed lines are Fermi surfaces of $f_{c,\mathbf{k}}$, $f_{v,\mathbf{k}+\mathbf{Q}}$ and $f_{c,\mathbf{k}-\mathbf{Q}}$. Momentum mesh of $79\times 79$.
The non-vanishing inter-layer superconducting instability is found to be a pair-density wave (PDW) state with pairing momentum $\mathbf{Q}$: $\langle f_{v,\mathbf{k}+\mathbf{Q}}f_{c,-\mathbf{k}}\rangle,\langle f_{c,\mathbf{k}+\mathbf{Q}}f_{v,-\mathbf{k}}\rangle\neq 0$ (as depicted in Fig. 5), as opposed to the standard zero-momentum pairing. As seen, the PDW superconductor is stabilized near the quantum critical line, where the interactions mediated by the soft bosons $b_{\pm}$ are strongest. The development of the PDW state requires a critical interaction strength, therefore, for small values of $g$ or large values of $m_{b}$, superconductivity disappears, as seen in Fig. 5. We find that the obtained PDW gap functions have an anisotropic $p$-wave character (see Fig. 5, bottom panel) with nodes at $k_{x}=0,-Q$. The resulting BdG spectrum consists of gapless quasiparticles that form a Fermi surface, as is typical for PDW superconductors [2].
@@
Experimental signatures.— The X-DW order as well as both the superconducting phases involve an inter-layer condensate (the Goldstone mediated superconductor also involves intra-layer pairing). This suggests drag transport – with separate electrical contacts, current is driven in one layer and the voltage generated is measured in the opposite layer – as a natural probe for these phenomena [18]. Analogous to the exciton condensates found in bilayer quantum Hall systems, the drag (and counterflow) resistivity will be zero in the X-DW phase even though the excitons condense at a finite momentum. We direct the reader to [1] for arguments based on effective response actions. Since the PDW superconductor is an interlayer Cooper pair condensate, the drag resistivity is zero even in this case. However, in contrast to the X-DW, the PDW state will exhibit superconductivity when the layers are contacted together. Finally, in the quantum critical NFL regime, the hotspot contribution to AC counterflow transport could potentially contain signatures of the quantum criticality, if they can be isolated from the contributions from the other “cold” parts of the Fermi surface [60].
For the predicted nodal superconductors, a direct signature is the characteristic ‘V-shape’ of the density of states from scanning tunneling microscopy (STM) [46, 28]. With the dual gate setup obstructing easy access the sample surface, a single-gate setup (that simultaneously tunes both density and bias potential) with an exposed top layer would allow a portion of the superconducting phase diagram to be directly imaged. In addition, measurements of the superfluid stiffness – via recently developed kinetic inductance techniques (see Refs. [5, 64]) – would also indicate the gapless BdG spectrum from its linear-in-temperature scaling behavior (in the clean limit [23, 49]).
@@
Outlook.— The two forms of superconductivity – one arising due to Goldstone fluctuations and another due to quantum critical fluctuations – can be realized at zero temperature. A natural direction to explore would be finite temperature energy competition between the two superconductors (for instance, the critical temperature $T_{c}$ of each and the dependence on $n,V_{b}$). It would also be interesting to incorporate dynamical screening effects to examine the magnitude of suppression of the exciton onset temperature, as compared to the zero-momentum exciton condensate [62].
@@
Recent experiments have presented evidence for the existence of a Fermi liquid state of trions (bound state of a fermion and an exciton) [42]. It would be interesting to incorporate trions in our framework (expected to become important for hole-electron densities of the ratio $p=2n$ [42]), which would provide an alternate route for the excitons to become gapped by starting from the exciton condensed phase (as studied in the cold-atom context in Ref. [3]).
@@
Finally, it would be intriguing to examine whether analogous unconventional superconductivity may be triggered in compensated twisted double bilayer graphene, where finite-momentum exciton condensates have been discussed [21, 20].
Acknowledgements
@@
Supplemental Material for “Unconventional superconductivity mediated by exciton density wave fluctuations”
I Microscopic Parameters of Model
In this study, we take WS2/WTe2 as a representative material in our numerics. This heterostructure is expected to have a type-II band alignment with the conduction (valence) band belonging to the WS2 (WTe2) layer. Let us define a microscopic lattice length scale $a=0.32\text{nm}$. This is not necessarily the lattice constant of the semiconductors under consideration, but a convenient length scale that we use to express other physical parameters. The effective valence and conduction masses are $m_{v,x}a^{2}=0.89\text{eV}^{-1}$, $m_{v,y}a^{2}=0.54\text{eV}^{-1}$[38], $m_{c,x}a^{2}=m_{c,y}a^{2}=0.33\text{eV}^{-1}$ [16], with a gate distance of $d_{g}=1.5\text{nm}$ and interlayer distance $d=3.5$ nm. We note that our parameter choice of $d_{g}$ is motivated by the MoSe22/WSe2 experimental study Ref. [42, 43], where the W and Mo layers are separated by a thin hBN spacer of thickness of 1.521.5-2nm in the channel region, while the exciton ‘contact’ region is separated by 102010-20 nm. For the Hubbard approximation models, we adopt a rectangular momentum grid of cutoff length $k_{c,x}=0.1a^{-1}$ along the $x$-direction, and $k_{c,y}=k_{c,x}\left(\frac{m_{v,y}m_{c,y}}{m_{v,x}m_{c,x}}\right)^{1/4}$ along the $y$-direction. The chemical potential and bias voltage defined in the main text determine the Fermi momenta $k_{F,c/v,x/y}$ of the carriers in the conduction/valence (cc/$v$) layers along the $\hat{x}/\hat{y}$ directions.
@@ I Microscopic Parameters of Model
II Density Wave Instability
At the non-interacting level, the system is described by an annualar Fermi surface, with an inner electron-like Fermi surface and an outer hole-like Fermi surface. Previous studies [50, 68] have established an instability of the annular Fermi surface to an interlayer charge density wave, which can be described by the order parameter $\langle f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}f_{c,\mathbf{k}}\rangle$. We obtain the conditions required for such an instability for our model within mean-field theory. Performing a mean-field decoupling in the spiral X-DW channel of interest, we obtain the following mean-field Hamiltonian:
@@ II Density Wave Instability
$\Sigma$ $=-\frac{1}{A}\sum_{\mathbf{k},\mathbf{k}^{\prime}}U(\mathbf{k}^{\prime}-\mathbf{k})f^{\dagger}_{c,\mathbf{k}}f_{v,\mathbf{k}+\mathbf{Q}}\langle f^{\dagger}_{v,\mathbf{k}^{\prime}+\mathbf{Q}}f_{c,\mathbf{k}^{\prime}}\rangle+h.c.$ (A.9)
@@
$H_{\text{mf}}=\sum_{\mathbf{k}}\begin{pmatrix}f^{\dagger}_{c,\mathbf{k}}&f^{\dagger}_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}\begin{pmatrix}\xi_{c,\mathbf{k}}&-\chi_{\mathbf{k}}\\ -\chi^{*}_{\mathbf{k}}&\xi_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}\begin{pmatrix}f_{c,\mathbf{k}}\\ f_{v,\mathbf{k}+\mathbf{Q}}\end{pmatrix}$ (A.10)
@@
$\chi_{\mathbf{k}}=\frac{1}{A}\sum_{\mathbf{k}^{\prime}}V(\mathbf{k}-\mathbf{k}^{\prime})\left(n_{F}(\xi_{\mathbf{k}^{\prime}}-E_{\mathbf{k}^{\prime}})-n_{F}(\xi_{\mathbf{k}^{\prime}}+E_{\mathbf{k}^{\prime}})\right)\frac{\chi_{\mathbf{k}^{\prime}}}{2E_{\mathbf{k}^{\prime}}}$ (A.11)
@@
This formulation to determine the exciton order parameter considers a select number of momentum shells/bands. For the exciton density wave instabilities (both spiral and collinear X-DW), we hence also perform fully self-consistent Hartree-Fock theory, including up to 7 $\times$2 bands; the number of bands selected is such that the UV cutoff is reached. We direct the reader to Ref. [17] for analogous Hartree-Fock calculations performed in rhombohedral stacked multilayer graphene. We choose a UV cutoff of $\pm 0.1/a$ in the $\hat{y}-$direction, and the first Brillouin zone along the $\hat{x}-$direction bounded by the $\pm|{\mathbf{Q}}|/2$. The dielectric constant is taken to be $\epsilon=10$; typical choices of the hBN dielectric environment is $5\sim 10$ to account for additional screening by the mobile charge carriers in a system with finite carrier density. As described in the main text, our focus is on the inter-layer instabilities, rather than the intra-layer instabilities. This amounts to suppressing the intra-layer Hartree and Fock contributions in the decoupling.
We present in Fig. A.1 a representative electronic bandstructure depicting the free (black) and finite exciton density wave order parameter (red). As seen, a electronic bands still cross the Fermi level, despite a finite exciton condensate, consistent with the three-band Hubbard approximation presented in Fig. 3 in Sec. III in the main text.
[…]
III Estimation of $V_{DW}$